OpenStudy (anonymous):

Solve the system by substitution 2.5x-3y=-13 3.25x-y=-14

OpenStudy (anonymous):

Use the second equation to make y the subject. Move y to the RHS while also moving -14 to the LHS. (LHS= Left Hand Side, RHS= Right Hand Side) \[y= 3.25x + 14\] In the first equation: Every time you see the pronumeral y, substitute it with (3.25x + 14) \[2.5x - 3(3.25x+14) = -13\] \[2.5x - 9.75x -42= -13\] Move the x components to the RHS and the numbers to the LHS. \[-42+13= 9.75x -2.5x\] \[29=7.25x\] \[x=\frac{ 29 }{ 7.25 }\] \[x= 4\]

OpenStudy (anonymous):

Thank you, again.

OpenStudy (anonymous):

No worries.